If you have a data set with n observations and want to construct a frequency distribution,...

  1. If you have a data set with n observations and want to construct a frequency distribution, what is a good approximation for the number of classes?
Answer #1

The approximate number of classes to construct a frequency distribution for n number of observations is obtain by using following formula.

K=1+3.322log10 n

where K= Number of classes

and log10 n = Logarithm of the total number of observations of base 10.

Example: If the total number of observations is 70, so the number of classes (K) are

K=1+3.322log10 100

K=1+3.322(2)

K= 1 + 6.644 = 7.644 = 8

K = 8

So for n = 100 we need approximate 8 classes.